# Random expansions of trees with bounded height

### v1

@@ 6. The base sequence of trees, expansions of them, and probabilities
We adopt the assumptions from the previous section, so $\sigma$ is a finite relational signature, $\tau\subseteq\sigma$, and $\tau=\{E\}$ where $E$ is a binary relation symbol. We consider a base sequence $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ where each $\mathcal{T}_{n}$ is a tree, so in particular $\mathcal{T}_{n}$ is a $\tau$-structure, and $|T_{n}|\to\infty$ as $n\to\infty$. Then $\mathbf{W}_{n}$ will be the set of all expansions of $\mathcal{T}_{n}$ to $\sigma$, and a probability distribution $\mathbb{P}_{n}$ will be defined on $\mathbf{W}_{n}$, via the concept of $PLA^{*}(\sigma)$-network. Our goal is to identify some constraints on $\mathbf{T}$ and on the $PLA^{*}(\sigma)$-network such that under these constraints we can prove results about asymptotic equivalence of complex formulas to simpler (closure basic) formulas, and derive convergence results.
For example, we could fix an integer $\Delta\geq 1$ and let $\mathcal{T}_{n}$ be a tree in which every element (that is, vertex) has at most $\Delta$ children. However, this case is covered by the context studied by Koponen in [20]. So we will allow, in fact require, that there is a function, say $g_{1}$, such that $\lim_{n\to\infty}g_{1}(n)=\infty$ and every nonleaf of $\mathcal{T}_{n}$ has at least $g_{1}(n)$ children. This condition alone does not imply the kind of results that we are looking for. So, in addition, we will assume that there is an integer $\Delta\geq 1$ such that, for all $n$, the height of $\mathcal{T}_{n}$ is at most $\Delta$. We justify this condition by the fact that such trees can represent hierarchies consisting of a fixed number of levels and where each “group” of objects in such a hierarchy has many more members than the number of levels. Moreover, such hierarchies are present in many situations and may influence relations which are modelled/expressed by probabilistic methods.
The perhaps simplest sequence $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ for which the main results of this article hold is obtained by fixing some integer $\Delta\geq 1$ and letting $\mathcal{T}_{n}$ be a tree of height $\Delta$ such that every leaf is on level $\Delta$ and all nonleaves have exactly $n$ children. But in order to increase the applicability of our results we wish prove results for more general sequences $\mathbf{T}$. For example, we want to allow different elements (that is, vertices) of $\mathcal{T}_{n}$ to have different numbers of children. However, if we put no constraints on the relative number of children that different elements of a tree can have, then we do not get the results that we aim for. The next couple of examples illustrate this and also motivate the constraints that we will impose on the sequence $\mathbf{T}$ of trees.
@@
Suppose that $n$ is a large and even. By Lemma 2.1 and since $|T_{n}|$ is bounded by a polynomial in $n$ we can argue (somewhat informally) like this. Let $a\in T_{n}$ be a child of the root such that such that $a$ is a child of the root with a child, hence $n$ children.
@@
Let $\sigma$ be as in Example 6.1. For odd $n\in\mathbb{N}^{+}$ let $\mathcal{T}_{n}$ be a tree of height $2$ such that the root has $n$ children, one child of the root has $2n^{3}$ children, and all other children of the root have $n$ children. For even $n\in\mathbb{N}^{+}$ let $\mathcal{T}_{n}$ be a tree of height $2$ such that the root has $n$ children, two children of the root has $n^{3}$ children, and all other children of the root have $n$ children. Let $\mathbf{W}_{n}$ be the set of all expansions of $\mathcal{T}_{n}$ to $\sigma$ and let $\mathbb{P}_{n}$ be the uniform probability distribution on $\mathbf{W}_{n}$.
@@
However, in general it is not sufficient to just have some (not too small) lower bound on the number of siblings in $\mathcal{T}_{n}$ with a similar number of children. But we actually need, for every fixed tree $\mathcal{T}^{\prime}$, to put a (not too large) lower bound on the number of siblings $a^{\prime}$ of (any vertex) $a$ such that $\mathsf{N}_{\mathcal{T}_{n}}(a^{\prime},\mathcal{T}^{\prime})$ is not too different from $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})$, where we recall that $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})$ is the number of subtrees of $\mathcal{T}_{n}$ that are rooted in $a$ and isomorphic to $\mathcal{T}^{\prime}$. (It is not hard to modify Example 6.2, by considering trees in which all leaves are on level 3, so that it shows that this stronger constraint is necessary.) Also, we will assume that there is a polynomial upper bound to the number of children that any member of $\mathcal{T}_{n}$ can have. This will allow us to use Corollary 2.2 in such a way that the main results follow.
@@
(ii) Suppose that $\mathcal{T}^{\prime}$ is a tree of heigt $\geq 1$. Let $a\in T_{n}$ and suppose that $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})>0$. Since every nonleaf of $\mathcal{T}_{n}$ has at least $g_{1}(n)$ children it follows that $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})\geq g_{1}(n)$ if $n$ is large enough. Moreover, if $n$ is large enough, then whether $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})>0$ or not depends only on which level $a$ belongs to, because all leaves of $\mathcal{T}_{n}$ are on level $\Delta$ and all nonleaves have at least $g_{1}(n)$ children where $g_{1}(n)\to\infty$.
@@
Observe that if $\sigma=\tau$ then the DAG of a $PLA^{*}(\sigma)$-network have an empty vertex set and hence there is no probability formula associated to it. Example 6.11 gives an example of a $PLA^{*}(\sigma)$-network.
@@ 4. $\theta_{R}(x,y)$ = “If $x$ is not on level 3 then 0, else
      • if $y$ belongs to the same subcommunity as $x$, then if $P_{3}(y)$ then $3/4$, else $1/2$, else
      • the proportion of $y$ on level 3 such that $P_{3}(y)$, divided by $3/4$, else
@@
    • if $y$ is on level 1 then the proportion of children of $y$ such that $P_{1}(y)$, divided by $3/2$.”
@@ 7.1. Definitions and some immediate consequences
Let $\varphi(\bar{x}),\psi(\bar{x})\in PLA^{*}(\sigma)$. It follows from the definition that if there are only finitely many $n$ such that there is $\bar{a}\in(T_{n})^{|\bar{x}|}$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\varphi(\psi(\bar{a})}\big{)}>0$ then, for every $\alpha$, $(\varphi,\psi)$ converges to $\alpha$ with respect to $\mathbb{G}$. It also follows from the definition that if there are infinitely many $n$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\varphi(\psi(\bar{a})}\big{)}>0$ for some $\bar{a}$, but there are only finitely many $n$ such that there is $\bar{a}\in(T_{n})^{|\bar{x}|}$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\varphi(\bar{a})\wedge\psi(\bar{a})}\big{)}>0$ then $(\varphi,\psi)$ converges to 0 with respect to $\mathbb{G}$.
@@
Let $\varphi(\bar{x},\bar{y}),\psi(\bar{x},\bar{y}),\chi(\bar{x})\in PLA^{*}(\sigma)$ be $0/1$-valued. Suppose that $\varphi\wedge\psi\wedge\chi$ is not cofinally satisfiable. We claim that then $(\varphi,\psi,\chi)$ is 0-balanced with respect to $\mathbb{G}$. It suffices to show that, for all sufficiently large $n$, if $\mathcal{A}\in\mathbf{W}_{n}$, $\mathcal{A}\models q(\bar{a})$, and $\mathcal{A}\models\psi(\bar{a},\bar{b})$, then $\mathcal{A}\not\models\varphi(\bar{a},\bar{b})$. But this is immediate from the assumption that $\varphi\wedge\psi\wedge\chi$ is not cofinally satisfiable. For this reason we may assume, when proving results about balanced triples, that the conjunction of the involved formulas is cofinally satisfiable and in particular consistent.
@@ 7.2. Proofs of results about convergence and balance
(2) Suppose that $p(\bar{x},\bar{y})$, $r(\bar{x},\bar{y})$ and $q(\bar{x})$ are closure types over $\tau$ and suppose that $p\wedge r\wedge q$ is cofinally satisfiable. Then $p$ and $r$ are equivalent and $p$ implies $q$. Suppose that $B\subseteq r(\bar{a},\mathcal{T}_{n})$, and $|B|\geq g_{3}(n)$. As $\lim_{n\to\infty}g_{3}(n)$ this means that $B\neq\emptyset$ if $n$ is large enough. Suppose that $\mathcal{A}\in\mathbf{W}_{n}$, so $\mathcal{A}=\mathcal{T}_{n}$. Also suppose that $\mathcal{T}_{n}\models q(\bar{a})$. Let $\bar{b}\in B$, so $\mathcal{T}_{n}\models r(\bar{a},\bar{b})$ and hence $\mathcal{T}_{n}\models p(\bar{a},\bar{b})$ (as $p$ and $r$ are equivalent). Thus $B\subseteq p(\bar{a},\mathcal{T}_{n})$ and $|B\cap p(\bar{a},\mathcal{A})|=|B|$, so for every $\varepsilon>0$, $(1-\varepsilon)|B|\leq|p(\bar{a},\mathcal{T}_{n})\cap B|\leq(1+\varepsilon)|B|.$ Since $\mathbb{P}_{n}(\mathcal{T}_{n})=1$ we get the conclusion of part (2). $\square$
@@
$\sigma_{\rho}=\{R\in\sigma:\text{ $R$ is on level $i$ of the underlying graph of $\mathbb{G}$ where $i\leq\rho$}\},$
@@ 7.3. Convergence
In this subsection we prove statement (B) above. We do it in three steps.
@@
Proof. As said in Remark 7.3 we may assume that there are infinitely many $n$ such that $p(\bar{x})\wedge p_{\tau}(\bar{x})$ is satisfiable in some $\mathcal{A}\in\mathbf{W}_{n}$, and it follows that $p\negthickspace\upharpoonright\negthickspace\tau$ is equivalent to $p_{\tau}$ (as every closure type over $\tau$ is a complete closure type over $\tau$, by the definition). Let $p_{\rho}=p\negthickspace\upharpoonright\negthickspace\sigma_{\rho}$, so $p_{\rho}$ is a complete closure type over $\sigma_{\rho}$. By Assumption 7.8 (2) there is $\alpha$ such that $(p_{\rho},p_{\tau})$ converges to $\alpha$ with respect to $\mathbb{G}$. By Lemma 7.11 there is $\beta$ such that $(p,p_{\rho})$ converges to $\beta$ with respect to $\mathbb{G}$. Hence, for all $\varepsilon>0$ there is $n_{0}$ such that for all $n\geq n_{0}$ and all $\bar{a}\in(T_{n})^{|\bar{x}|}$,
@@
By Corollary 7.13, part (2) of Assumption 7.8 holds if $\sigma_{\rho}$ and $\mathbb{G}_{\rho}$ are replaced by $\sigma$ and $\mathbb{G}$, respectively. So the induction step for convergence is completed, that is, claim (B) above is proved.

### v2

@@ 6. The base sequence of trees, expansions of them, and probabilities
We adopt the assumptions from the previous section, so $\sigma$ is a finite relational signature, $\tau\subseteq\sigma$, and $\tau=\{E\}$ where $E$ is a binary relation symbol. We consider a base sequence $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ where each $\mathcal{T}_{n}$ is a tree, so in particular $\mathcal{T}_{n}$ is a $\tau$-structure, and $|T_{n}|\to\infty$ as $n\to\infty$. Then $\mathbf{W}_{n}$ will be the set of all expansions of $\mathcal{T}_{n}$ to $\sigma$, and a probability distribution $\mathbb{P}_{n}$ will be defined on $\mathbf{W}_{n}$ via a $PLA^{*}(\sigma)$-network. Our goal is to identify some constraints on $\mathbf{T}$ and on the $PLA^{*}(\sigma)$-network such that under these constraints we can prove results about asymptotic equivalence of complex formulas to simpler (closure basic) formulas, and derive convergence results.
For example, we could fix an integer $\Delta\geq 1$ and let $\mathcal{T}_{n}$ be a tree in which every element (that is, vertex) has at most $\Delta$ children. However, this case is covered by the context studied by Koponen in [20]. So we will allow, in fact require, that there is a function, say $g_{1}$, such that $\lim_{n\to\infty}g_{1}(n)=\infty$ and every nonleaf of $\mathcal{T}_{n}$ has at least $g_{1}(n)$ children. This condition alone does not imply the kind of results that we are looking for. So, in addition, we will assume that there is an integer $\Delta\geq 1$ such that, for all $n$, the height of $\mathcal{T}_{n}$ is at most $\Delta$.
The perhaps simplest sequence $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ for which the main results of this article hold is obtained by fixing some integer $\Delta\geq 1$ and letting $\mathcal{T}_{n}$ be a tree of height $\Delta$ such that every leaf is on level $\Delta$ and all nonleaves have exactly $n$ children. But in order to increase the applicability, we prove results for more general sequences $\mathbf{T}$. For example, we want to allow different elements (that is, vertices) of $\mathcal{T}_{n}$ to have different numbers of children. However, if we put no constraints on the relative number of children that different elements of a tree can have, then we do not get the results that we aim for. The next couple of examples illustrate this and also motivate the constraints that we will impose on the sequence $\mathbf{T}$ of trees.
@@
Suppose that $n$ is a large and even. By Lemma 2.1 and since $|T_{n}|$ is bounded by a polynomial in $n$ we can argue (somewhat informally) like this. Let $a\in T_{n}$ be a child of the root such that $a$ is a child of the root with a child, hence $n$ children.
@@
Let $\sigma$ be as in Example 6.1. For odd $n\in\mathbb{N}^{+}$ let $\mathcal{T}_{n}$ be a tree of height $2$ such that the root has $n$ children, one child of the root has $2n^{3}$ children, and all other children of the root have $n$ children. For even $n\in\mathbb{N}^{+}$ let $\mathcal{T}_{n}$ be a tree of height $2$ such that the root has $n$ children, two children of the root have $n^{3}$ children, and all other children of the root have $n$ children. Let $\mathbf{W}_{n}$ be the set of all expansions of $\mathcal{T}_{n}$ to $\sigma$ and let $\mathbb{P}_{n}$ be the uniform probability distribution on $\mathbf{W}_{n}$.
@@
However, in general it is not sufficient to just have some (not too small) lower bound on the number of siblings in $\mathcal{T}_{n}$ with a similar number of children. But we actually need, for every fixed tree $\mathcal{T}^{\prime}$, to put a (not too large) lower bound on the number of siblings $a^{\prime}$ of (any vertex) $a$ such that $\mathsf{N}_{\mathcal{T}_{n}}(a^{\prime},\mathcal{T}^{\prime})$ is not too different from $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})$, where we recall that $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})$ is the number of subtrees of $\mathcal{T}_{n}$ that are rooted in $a$ and isomorphic to $\mathcal{T}^{\prime}$. (It is not hard to modify Example 6.2, by considering trees in which all leaves are on level 3, so that it shows that this stronger constraint is necessary.) Also, we will assume that there is a polynomial upper bound on the number of children that any member of $\mathcal{T}_{n}$ can have. This will allow us to use Corollary 2.2 in such a way that the main results follow.
@@
Although the above assumption may look complicated, the requirement on $g_{1},g_{2}$ and $g_{3}$ is just that they grow faster than every logarithm and slower that some polynomial, and that $g_{1}$ grows faster than both $g_{2}$ and $g_{3}$. So there are many possible choices of such functions, for example $g_{1}(n)=(\ln n)^{3}$, $g_{2}(n)=\ln n$, and $g_{3}(n)=(\ln n)^{2}$ for all $n>1$.
@@
(ii) Suppose that $\mathcal{T}^{\prime}$ is a tree of height $\geq 1$. Let $a\in T_{n}$. For all sufficiently large $n$ we have that if $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})>0$, then $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})\geq g_{1}(n)$ as we now demonstrate. Suppose that $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})>0$. Then there is a subtree $\mathcal{T}^{\prime\prime}$ of $\mathcal{T}_{n}$ that is isomorphic to $\mathcal{T}^{\prime}$ and rooted in $a$. Let $b$ be a nonleaf vertex of $\mathcal{T}^{\prime\prime}$ such that all children of $b$ are leaves. Let $b_{1},\ldots,b_{m}$ be all children of $b$ in $\mathcal{T}_{n}$, so $m\geq g_{1}(n)$, and let $b^{\prime}_{1},\ldots,b^{\prime}_{k}$ be all children of $b$ in $\mathcal{T}^{\prime\prime}$. Assuming that $n$ is large enough, $\{b^{\prime}_{1},\ldots,b^{\prime}_{k}\}$ is a proper subset of $\{b_{1},\ldots,b_{m}\}$. For every choice of $B\subseteq\{b_{1},\ldots,b_{m}\}$ such that $|B|=k$ we get a subtree $\mathcal{T}_{B}$ of $\mathcal{T}_{n}$ that is rooted in $a$ and isomorphic to $\mathcal{T}^{\prime\prime}$ (hence to $\mathcal{T}^{\prime}$) by replacing $b^{\prime}_{1},\ldots,b^{\prime}_{k}$ in $\mathcal{T}^{\prime\prime}$ by the vertices in $B$. The subtree $\mathcal{T}_{B}$ can also be described as the subtree of $\mathcal{T}_{n}$ generated by $(T^{\prime\prime}\setminus\{b^{\prime}_{1},\ldots,b^{\prime}_{k}\})\cup B$. There are $\binom{m}{k}\geq\binom{g_{1}(n)}{k}\geq g_{1}(n)$ choices of $B$ and therefore $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})\geq g_{1}(n)$.
Moreover, if $n$ is large enough, then whether $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})>0$ or not depends only on which level $a$ belongs to, because all leaves of $\mathcal{T}_{n}$ are on level $\Delta$ and all nonleaves have at least $g_{1}(n)$ children where $g_{1}(n)\to\infty$.
@@
Observe that if $\sigma=\tau$ then the DAG of a $PLA^{*}(\sigma)$-network has an empty vertex set and hence there is no probability formula associated to it. Example 6.11 gives an example of a $PLA^{*}(\sigma)$-network.
@@ 4. $\theta_{R}(x,y)$ = “If $x$ is not on level 3 then 0, else
      • if $y$ is a sibling of $x$, then if $P_{3}(y)$ then $3/4$, else $1/2$, else
      • the proportion of $y$ on level 3 such that $P_{3}(y)$, multiplied by $3/4$, else
@@
    • if $y$ is on level 1 then the proportion of children of $y$ such that $P_{1}(y)$, multiplied by $2/3$.”
@@ 7.1. Definitions and some immediate consequences
Let $\varphi(\bar{x}),\psi(\bar{x})\in PLA^{*}(\sigma)$. It follows from the definition that if there are only finitely many $n$ such that there is $\bar{a}\in(T_{n})^{|\bar{x}|}$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\psi(\bar{a})}\big{)}>0$ then, for every $\alpha$, $(\varphi,\psi)$ converges to $\alpha$ with respect to $\mathbb{G}$. It also follows from the definition that if there are infinitely many $n$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\psi(\bar{a})}\big{)}>0$ for some $\bar{a}$, but there are only finitely many $n$ such that there is $\bar{a}\in(T_{n})^{|\bar{x}|}$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\varphi(\bar{a})\wedge\psi(\bar{a})}\big{)}>0$ then $(\varphi,\psi)$ converges to 0 with respect to $\mathbb{G}$.
@@
Let $\varphi(\bar{x},\bar{y}),\psi(\bar{x},\bar{y}),\chi(\bar{x})\in PLA^{*}(\sigma)$ be $0/1$-valued. Suppose that $\varphi\wedge\psi\wedge\chi$ is not cofinally satisfiable. We claim that then $(\varphi,\psi,\chi)$ is 0-balanced with respect to $\mathbb{G}$. It suffices to show that, for all sufficiently large $n$, if $\mathcal{A}\in\mathbf{W}_{n}$, $\mathcal{A}\models\chi(\bar{a})$, and $\mathcal{A}\models\psi(\bar{a},\bar{b})$, then $\mathcal{A}\not\models\varphi(\bar{a},\bar{b})$. But this is immediate from the assumption that $\varphi\wedge\psi\wedge\chi$ is not cofinally satisfiable. For this reason we may assume, when proving results about balanced triples, that the conjunction of the involved formulas is cofinally satisfiable and in particular consistent.
@@ 7.2. Proofs of results about convergence and balance
(2) Suppose that $p(\bar{x},\bar{y})$, $r(\bar{x},\bar{y})$ and $q(\bar{x})$ are closure types over $\tau$ and suppose that $p\wedge r\wedge q$ is cofinally satisfiable. Then $p$ and $r$ are equivalent and $p$ implies $q$. Suppose that $B\subseteq r(\bar{a},\mathcal{T}_{n})$, and $|B|\geq g_{3}(n)$. As $\lim_{n\to\infty}g_{3}(n)=\infty$ this means that $B\neq\emptyset$ if $n$ is large enough. Suppose that $\mathcal{A}\in\mathbf{W}_{n}$, so $\mathcal{A}=\mathcal{T}_{n}$. Also suppose that $\mathcal{T}_{n}\models q(\bar{a})$. Let $\bar{b}\in B$, so $\mathcal{T}_{n}\models r(\bar{a},\bar{b})$ and hence $\mathcal{T}_{n}\models p(\bar{a},\bar{b})$ (as $p$ and $r$ are equivalent). Thus $B\subseteq p(\bar{a},\mathcal{T}_{n})$ and $|B\cap p(\bar{a},\mathcal{A})|=|B|$, so for every $\varepsilon>0$, $(1-\varepsilon)|B|\leq|p(\bar{a},\mathcal{T}_{n})\cap B|\leq(1+\varepsilon)|B|.$ Since $\mathbb{P}_{n}(\mathcal{T}_{n})=1$ we get the conclusion of part (2). $\square$
@@
$\sigma_{\rho}=\tau\cup\{R\in\sigma\setminus\tau:\text{ $R$ is on level $i$ of the underlying graph of $\mathbb{G}$ where $i\leq\rho$}\},$
@@ 7.3. Convergence
In this subsection we prove statement (A) above. We do it in three steps.
@@
Proof. As said in Remark 7.3 we may assume that there are infinitely many $n$ such that $p(\bar{x})\wedge p_{\tau}(\bar{x})$ is satisfiable in some $\mathcal{A}\in\mathbf{W}_{n}$, and it follows that $p\negthickspace\upharpoonright\negthickspace\tau$ is equivalent to $p_{\tau}$ (as every closure type over $\tau$ is a complete closure type over $\tau$, by the definition). Let $p_{\rho}=p\negthickspace\upharpoonright\negthickspace\sigma_{\rho}$, so $p_{\rho}$ is a complete closure type over $\sigma_{\rho}$. By Assumption 7.8 (2) there is $\alpha$ such that $(p_{\rho},p_{\tau})$ converges to $\alpha$ with respect to $\mathbb{G}_{\rho}$. By Lemma 7.11 there is $\beta$ such that $(p,p_{\rho})$ converges to $\beta$ with respect to $\mathbb{G}$. Hence, for all $\varepsilon>0$ there is $n_{0}$ such that for all $n\geq n_{0}$ and all $\bar{a}\in(T_{n})^{|\bar{x}|}$,
@@
By Corollary 7.13, part (2) of Assumption 7.8 holds if $\sigma_{\rho}$ and $\mathbb{G}_{\rho}$ are replaced by $\sigma$ and $\mathbb{G}$, respectively. So the induction step for convergence is completed, that is, claim (A) above is proved.